The change of internal energy of an ideal gas is 300 J. What is the total heat added into the gas if it expands for 2.010−3 m3 under 1.01105 Pa
1
Expert's answer
2019-12-05T10:23:05-0500

1st law of thermodynamics give it


"Q=\Delta U+p\Delta V"

Initial conditions


"\Delta U=300; p=1.01 \cdot 10^5 (Pa); \Delta V=2.0\cdot 10^{-3} (m^3)"

Calculating


"Q=300+1.01\cdot 10^5 \cdot 2.0\cdot 10^{-3}=502 (J)"